Date: 10/21/18 incubation times are approximately normally distributed with a standard deviation of 1 day. Complete parts (a) through (c) below Click here to view.page 1 of the standard normal distribution table, Click here to view page 2 of the standard normal distribution table 4. The mean incubation time for a type of fertilized egg kept at 100.8 F is 23 days. Suppose that the (a) What is the probability that a randomly selected fertilized egg hatches in shorter than 21 days? The probability that a randomly selected fertilized egg hatches in shorter than 21 days is (Round to four decimal places as needed.) (tb) What is the probablity that a randomly selected fertilized egg hatches between 22 and 23 days? The probability that a randomly selected fertilized egg hatches between 22 and 23 days is (Round to four decimal places as needed.) (c) What is the probability that a randomly selected fertilized egg takes longer than 25 days to hatch? The probability that a randomly selected fertilized egg takes longer than 25 days to hatch is Round to four decimal places as needed.) 1: Standard Normal Distribution Table (Page 1) 00.8°F below table able, egg ha s in sh .n0
Answer #1

a. P(X<21) = P(Z<21-23/1)

=P(Z<-2) =============>use excel formula ''=NORMSDIST(-2)

=0.02275 =0.02278 (In four decimal)

b)P(22<X<23)=P(X<23)-P(X<22)

=P(Z<23-23/1)-P(Z<22-23/1)

=P(Z<0)-P(Z<-1)

=0.5-01587

=0.34134=0.3413

c.P(X>25)=1-P(X<25)

=1-P(Z<25-23/1)

=1-P(Z<2)

=1-0.97725

=0.02275 =0.02278   

Hope this will be helpful. Thanks and God Bless You :)